# Growth of Light-Seed Black Holes in Gas-rich Galaxies at High Redshift

### v1

@@ Abstract
Recent observations by the James Webb Space Telescope confirm the existence of massive black holes ($>10^{6}$ $\ \text{M}_{\odot}$) beyond the redshift of $z=10$. However, their formation mechanism(s) still remain an open question. Light seed black holes are one such formation pathway, forming as the end stage of metalfree (Population III) stars. Light seed black holes can grow into massive black holes as long as they accrete near the Eddington limit for substantial periods or undergo several bursts of super-Eddington accretion. In this work, our aim is to ascertain if light seeds can grow in gas rich galaxies - similar to those expected at high redshift (z $\gtrsim 10$). Using the Arepo code, we follow self-consistently the formation of Population III stars and black holes in galaxies with total masses in the range $10^{8}$ $\rm{M_{\odot}}$. We find that in the absence of feedback, black holes can grow to $10^{5}$ $\ \text{M}_{\odot}$ in just $10^{4}$ years. These black holes do not decouple from the gas clumps in which they are born and are able to accrete at hyper-Eddington rates. In the presence of supernova feedback, the number of actively growing black holes diminishes by an order of magnitude. However, we still observe hyper-Eddington accretion in approximately 1% of the black hole population despite supernova feedback. This (idealised) work lays the foundation for future works, where we will test our models in a cosmological framework.
@@ 1 Introduction
Quasars at redshifts, $z\sim$ $6-7$ have been observed to host (super-)massive black holes (MBHs) [^1] We use the term MBH here to refer to central or near-central black holes with masses in excess of $10^{3}$ $\rm{M_{\odot}}$ of mass $10^{8-10}$ $\ \text{M}_{\odot}$ (Banados et al., 2018; Farina et al., 2022; Mazzucchelli et al., 2023; Wang et al., 2021; Yang et al., 2020, 2023) (see also review by Fan et al. 2023). Moreover, recent JWST observations at redshift of $z\sim 8-11$ have unearthed galaxies with central MBHs with masses of $\gtrsim 10^{6}$ $\ \text{M}_{\odot}$ (Maiolino et al., 2024; Larson et al., 2023; Goulding et al., 2023; Bogdán et al., 2024). These observations imply that MBHs must have grown rapidly from their seed masses at very early epochs. The physical origin of the seeds is still uncertain with competing models including remnants of Population III (PopIII) stars (Light Seeds - see e.g. Madau & Rees 2001), run-away mergers of stellar bodies in dense star clusters (e.g. Devecchi & Volonteri, 2009), or direct collapse of supermassive stars of $10^{4}-10^{6}$ $\ \text{M}_{\odot}$ (Heavy Seeds, e.g. Haehnelt & Rees 1993; Ferrara et al. 2014; Mayer & Bonoli 2018; Woods et al. 2017, 2019a). For this paper we define light seeds as those with (initial) masses of less than 1000 $\rm{M_{\odot}}~{}$and heavy seeds as those with masses in excess of 1000 $\rm{M_{\odot}}$.
@@
Lights seeds are expected to be formed at early times, starting anywhere from as high a redshift as $z\sim 30$ (Haiman & Loeb, 2001; Madau & Rees, 2001; Heger et al., 2003; Volonteri et al., 2003; Madau et al., 2004; Volonteri, 2012; Schauer et al., 2021) in low-mass dark matter halos of $\gtrsim 10^{6}$ $\ \text{M}_{\odot}$. Their progenitor (Pop III) stars exhibit mass ranges from tens to hundreds of solar masses (Bromm et al., 1999; Abel et al., 2000, 2002; Bromm et al., 2002; O’Shea & Norman, 2007; Turk et al., 2009; Clark et al., 2011; Hirano et al., 2014; Prole et al., 2023). Light seeds form through a multitude of pathways owing to the wide range of Pop III stellar masses including core-collapse supernovae (SNe) or hypernovae ($11\ \text{M}_{\odot}\leq M\leq 40\ \text{M}_{\odot}$, Woosley & Weaver 1995; Nomoto et al. 2006) and direct collapse into stellar mass black holes (BH) [^2] We use the abbreviation BH to refer implicitly to stellar mass black holes or light seeds ($40\ \text{M}_{\odot}\leq M\leq 140\ \text{M}_{\odot}$ and $M>260\ \text{M}_{\odot}$, Heger & Woosley 2002). These seeds can grow into a MBH if they have sustained near-Eddington accretion on Gyr time periods. However, this condition is non-trivial to satisfy in dark matter halos as feedback from progenitor stars, SNe explosion, BH accretion and BH dynamics affects the supply of gas to the BHs (Wise et al., 2008; Alvarez et al., 2009; Milosavljević et al., 2009; Smith et al., 2018), resulting in highly sub-Eddington accretion for long periods. The growth can be accelerated if the BHs can accrete at hyper-Eddington rates ($\rm{f_{Edd}}\gtrsim 1000$, see theoretical arguments by Inayoshi et al. 2016; Jiang et al. 2019; Park et al. 2020; Kitaki et al. 2021; Botella et al. 2022). However, these theoretical studies assume that the BH is able to pull gas from scales larger than their accretion disk or equivalently that gas flows rapidly onto the BHs from large distances. It is not clear if the BH will encounter dense regions of gas for such high accretion rates to be realised. If the BH does encounter such regions, it can grow into a MBH on timescales of less than 500 Myr (Lupi et al., 2016), or perhaps even less due to a rapid runaway accretion episode (Shi et al., 2023).
@@
More exotic, or equivalently rarer, channels through which MBHs seeds can be formed can avoid the need for super or hyper-Eddington accretion by forming with initially much heavier masses - so-called heavy seeds (Regan & Volonteri, 2024). As defined above, heavy seeds are seeds with masses in excess of 1000 $\rm{M_{\odot}}$. These can form from a number of potential channels including runaway collapse inside a dense stellar cluster (Portegies Zwart et al., 2004; Devecchi & Volonteri, 2009; Katz et al., 2015; Rizzuto et al., 2021; Arca Sedda et al., 2021; González et al., 2021; Reinoso et al., 2023; Arca Sedda et al., 2023; Rantala et al., 2024) or through runaway collisions inside a BH cluster (Fragione & Leigh, 2018; Fragione et al., 2022; Antonini et al., 2019; Mapelli et al., 2021; Arca Sedda et al., 2021). Both of these avenues have been extensively modelled over the previous two decades and while a full consensus is still lacking, forming MBH with masses in excess of $\sim 1000$ $\rm{M_{\odot}}~{}$appears viable.
@@
Another route to heavy seeds is via the intermediate stage of a very massive or supermassive PopIII star (Regan & Haehnelt, 2009a, b; Regan et al., 2014; Spaans & Silk, 2006; Latif et al., 2013a, b; Becerra et al., 2015; Chon et al., 2016; Chon & Latif, 2017; Chon et al., 2018; Chon & Omukai, 2020; Woods et al., 2019b, 2017, a; Wise et al., 2019; Regan et al., 2020). In this case a star with a mass between $10^{3-5}$ $\rm{M_{\odot}}~{}$forms and directly collapses to an near-equivalent mass black hole. Such heavy seeds are regularly invoked to explain the existence of MBH at high-z, while the chemical signatures of the progenitor super massive star (SMS) phase may already have been detected (Nandal et al., 2024a, b; Schaerer et al., 2024). However, in this study we will focus on light seeds and how rapid growth onto light seeds can provide an additional channel to MBH seeding.
@@
To understand the potential hurdles to the growth of light seeds, we study here whether rapid growth of light seeds can be achieved within gas-rich galaxies at high redshift. We conducted simulations using the Arepo code (Springel, 2010; Pakmor et al., 2016), incorporating self-consistent PopIII star formation, MBH formation as well as SNe feedback. Our newly developed modules are designed to simulate the creation and accretion of sink particles, which serve as analogs to stars during their stellar lifetimes and subsequently evolve into BHs.
@@
The paper is structured as follows. In §2, we discuss the numerical methodology our work. We briefly describe the working of Arepo code. We also discuss in depth the initial conditions used and the sink particle techniques employed. §3 describes our results for two realisations (with and without SNe feedback). We discuss some of our caveats and compare our study with others from the literature in §4 and finally conclude our work in §5.
2 Numerical method
To carry out the simulations described in this paper, we use the moving mesh code Arepo (Springel, 2010; Pakmor et al., 2016), which solves the Euler equations on an unstructured Voronoi tessellation that moves with the flow. The fluid dynamics are solved using a finite-volume, second-order reconstruction scheme coupled with an exact Riemann solver. The unstructured mesh adapts by locally refining and de-refining in a quasi-Lagrangian way to maintain a typical gas mass target per cell. We employ the hierarchical time integration scheme described in Springel et al. (2021), which splits the Hamiltonian into “fast” and “slow” components, i.e. those on short versus long time-steps, respectively. With this approach, interactions between particles/cells on the smallest time steps (up to maximum of 3000 particles) are solved using direct summation, while a standard octree approach is used for all other interactions. This is primarily done for efficiency, to avoid unnecessary tree constructions for small numbers of particles, but has the added benefit of more accurately calculating the gravitational forces between the black holes and their nearest neighbours, which are on the shortest timesteps, via direct summation (Bourne et al., 2023).
Within our suite of simulations, the dark matter (DM) particles have gravitational softening lengths $\Delta x$ = $10^{-2}$ pc, while gas cell softening adapt with the cell size down to a minimum of $\Delta x$ = $10^{-2}$ pc. This is the minimum cell length we allow and is controlled by the MinVolume parameter. Refinement is controlled via the JeansRefinement parameter with the number of cells used to resolve the Jeans length of the gas set to 8.
@@ 2.1 Initial conditions
We create the initial conditions (ICs) for our idealised and isolated galaxies using GalIC (Yurin & Springel, 2014), which constructs galaxy models in collisionless equilibrium made up of both DM and other collisionless particles which make up the disk and bulge components. To accurately and realistically model high redshift early galaxies, we initialise our galaxies using the results from Chen et al. 2014, who found that for halo masses between $10^{7}$ $\rm{M_{\odot}}~{}$and $10^{8.5}$ $\rm{M_{\odot}}~{}$, the gas fraction to the total galaxy mass is $0.1$, lower than the cosmic baryon fraction. The reason for this is that in small, early, galaxies SNe feedback is very effective in removing gas from the galaxies, at least, until galaxies become more massive. As a result, we create our idealised galaxies with $90\%$ of the galaxy’s mass in the surrounding DM halo. The remaining mass of the galaxy being baryonic.
The DM halo is modeled as a NFW (Navarro et al., 1997) halo profile, while the gas is modeled as a spherical bulge with radius $0.1$ times the parent halo radius. We set the virialized circular velocity of the halo to be $10$ km/s, which constrains the total mass of the halo to be $2.32\times 10^{8}$ $\ \text{M}_{\odot}$. We create the ICs using $10^{6}$ DM particles and $10^{4}$ gas particles. The (initial) mass resolution is then $\sim 200$ $\rm{M_{\odot}}~{}$and $\sim 2000$ $\ \text{M}_{\odot}$ for the DM and gas particles, respectively. During the course of the simulation, the mass resolution of the gas cells refine based on local gas properties (specifically the Jeans length) and more gas cells are added to the grid as discussed in §2. The DM particle resolution remains fixed throughout the simulation but at $\sim 200$ $\rm{M_{\odot}}$, it is highly resolved. In addition to the DM particles and gas particles we also include $10^{4}$ collisionless ‘star’ particles which reside in a plane around the halo centre. These "stars" model the previous generation of stars inherited via the hierarchical build up of the galactic structure. The total mass in these genetic stars is $2.32\times 10^{6}$ and they add extra potential to the halo centre helping to stabilise the inner part of the galaxy. The different constituents to the halo are outlined in Table 1.
@@
Following the creation of the initial conditions using GalIC, the output must be reformatted into an output suitable for Arepo . To enable this, the ICs generated by GalIC must be inserted into our 20 kpc box and evolved using Arepo , this requires that the smoothed particle hydrodynamic (SPH) particles, created by GalIC, must be converted into mesh cells. This is achieved using the ADD_BACKGROUND functionality (Springel, 2010; Weinberger et al., 2020) in Arepo . This procedure of taking the idealised halo and inserting it into an Arepo box inevitably results in some mass loss from the halo. This is because a small number of particles are scattered far from the host halo due to dynamical interactions at the outskirts of the halo. These particles therefore leave the 20 kpc box (isolated boundary conditions) in which we conduct the Arepo simulation. The final result is therefore a halo of mass $1.78\times 10^{8}$ $\ \text{M}_{\odot}$ in a box of $20$ kpc.

### v2

@@ Abstract
Recent observations by the James Webb Space Telescope confirm the existence of massive black holes ($>10^{6}$ $\ \text{M}_{\odot}$) beyond the redshift of $z=10$. However, their formation mechanism(s) still remains an open question. Light seed black holes are one such formation pathway, forming as the end stage of metal-free (Population III) stars. Light seed black holes can grow into massive black holes as long as they accrete near the Eddington limit for substantial periods or undergo several bursts of super-Eddington accretion. In this work, our aim is to ascertain if light seeds can grow in gas-rich galaxies - similar to those expected at high redshift (z $\gtrsim 10$). Using the Arepo code, we follow self-consistently the formation of Population III stars and black holes in galaxies with total masses in the range $10^{8}$ $\rm{M_{\odot}}$. We find that in the absence of feedback, black holes can grow to $10^{5}$ $\ \text{M}_{\odot}$ in just $10^{4}$ yr. These black holes do not decouple from the gas clumps in which they are born and are able to accrete at hyper-Eddington rates. In the presence of supernova feedback, the number of actively growing black holes diminishes by an order of magnitude. However, we still observe hyper-Eddington accretion in approximately 1% of the black hole population despite supernova feedback. This (idealised) work lays the foundation for future works, where we will test our models in a cosmological framework. In this work, we neglect radiative feedback processes from stellar evolution and from accretion onto the growing black holes. This likely means that our results represent an upper limit to light seed growth. We will address these shortcomings in future work.
@@ 1 Introduction
Quasars at redshifts, $z\sim$ $6-7$ have been observed to host (super-)massive black holes (MBHs) [^1] We use the term MBH here to refer to central or near-central black holes with masses in excess of $10^{3}$ $\rm{M_{\odot}}$ of mass $10^{8-10}$ $\ \text{M}_{\odot}$ (Banados et al., 2018; Farina et al., 2022; Mazzucchelli et al., 2023; Wang et al., 2021; Yang et al., 2020, 2023) (see also review by Fan et al. 2023). Moreover, recent JWST observations at redshift of $z\sim 8-11$ have unearthed galaxies with central MBHs with masses of $\gtrsim 10^{6}$ $\ \text{M}_{\odot}$ (Maiolino et al., 2024; Larson et al., 2023; Goulding et al., 2023; Bogdán et al., 2024). These observations imply that MBHs must have grown rapidly from their seed masses at very early epochs. The physical origin of the seeds is still uncertain with competing models including remnants of Population III (PopIII) stars (Light Seeds - see e.g. Madau & Rees 2001), run-away mergers of stellar bodies in dense star clusters (e.g. Devecchi & Volonteri, 2009), or direct collapse of supermassive stars of $10^{4}-10^{6}$ $\ \text{M}_{\odot}$ (Heavy Seeds, see e.g. Haehnelt & Rees 1993; Ferrara et al. 2014; Mayer & Bonoli 2018; Woods et al. 2017, 2019a). For this paper, we define light seeds as those with (initial) masses of less than 1000 $\rm{M_{\odot}}~{}$and heavy seeds as those with masses in excess of 1000 $\rm{M_{\odot}}$.
@@
Lights seeds are expected to be formed at early times, starting anywhere from as high a redshift as $z\sim 30$ (Haiman & Loeb, 2001; Madau & Rees, 2001; Heger et al., 2003; Volonteri et al., 2003; Madau et al., 2004; Volonteri, 2012; Schauer et al., 2021) in low-mass dark matter halos of $\gtrsim 10^{6}$ $\ \text{M}_{\odot}$. These early halos will eventually merge to become some of the most massive halos in existence today. Their progenitor (Pop III) stars exhibit mass ranges from tens to hundreds of solar masses (Bromm et al., 1999; Abel et al., 2000, 2002; Bromm et al., 2002; O’Shea & Norman, 2007; Turk et al., 2009; Clark et al., 2011; Hirano et al., 2014; Prole et al., 2023). Light seeds form through a multitude of pathways owing to the wide range of Pop III stellar masses including core-collapse supernovae (SNe) or hypernovae ($11\ \text{M}_{\odot}\leq M\leq 40\ \text{M}_{\odot}$, Woosley & Weaver 1995; Nomoto et al. 2006) and direct collapse into stellar mass black holes (BH) [^2] We use the abbreviation BH to refer implicitly to stellar mass black holes or light seeds ($40\ \text{M}_{\odot}\leq M\leq 140\ \text{M}_{\odot}$ and $M>260\ \text{M}_{\odot}$, Heger & Woosley 2002). These seeds can grow into a MBH if they have sustained near-Eddington accretion on Gyr time periods. However, this condition is non-trivial to satisfy in dark matter halos as feedback from progenitor stars, SNe explosion, BH accretion and BH dynamics affects the supply of gas to the BHs (Wise et al., 2008; Alvarez et al., 2009; Milosavljević et al., 2009; Smith et al., 2018), resulting in highly sub-Eddington accretion for long periods. The growth can be accelerated if the BHs can accrete at hyper-Eddington rates ($\rm{f_{Edd}}\gtrsim 1000$, see theoretical arguments by Inayoshi et al. 2016; Jiang et al. 2019; Park et al. 2020; Kitaki et al. 2021; Botella et al. 2022). However, these theoretical studies assume that the BH is able to pull gas from scales larger than their accretion disk, or equivalently that gas flows rapidly onto the BHs from large distances. It is not clear if the BH will encounter dense regions of gas for such high accretion rates to be realised. If the BH does encounter such regions, it can grow into a MBH on timescales of less than 500 Myr (Lupi et al., 2016), or perhaps even less due to a rapid runaway accretion episode (Shi et al., 2023).
@@
More exotic, or equivalently rarer, channels through which MBH seeds can be formed can avoid the need for super- or hyper-Eddington accretion by forming with initially much heavier masses - so-called heavy seeds (Regan & Volonteri, 2024). As defined above, heavy seeds are seeds with masses in excess of 1000 $\rm{M_{\odot}}$. These can form from a number of potential channels, including runaway collapse inside a dense stellar cluster (Portegies Zwart et al., 2004; Devecchi & Volonteri, 2009; Katz et al., 2015; Rizzuto et al., 2021; Arca Sedda et al., 2021; González et al., 2021; Reinoso et al., 2023; Arca Sedda et al., 2023; Rantala et al., 2024) or through runaway collisions inside a BH cluster (Fragione & Leigh, 2018; Fragione et al., 2022; Antonini et al., 2019; Mapelli et al., 2021; Arca Sedda et al., 2021). Both of these avenues have been extensively modelled over the previous two decades and while a full consensus is still lacking, the formation of MBH with masses in excess of $\sim 1000$ $\rm{M_{\odot}}~{}$appears viable.
@@
Another route to heavy seeds is via the intermediate stage of a very massive or supermassive PopIII star (Regan & Haehnelt, 2009a, b; Regan et al., 2014; Spaans & Silk, 2006; Latif et al., 2013a, b; Becerra et al., 2015; Chon et al., 2016; Chon & Latif, 2017; Chon et al., 2018; Chon & Omukai, 2020; Woods et al., 2019b, 2017, a; Wise et al., 2019; Regan et al., 2020). In this case a star with a mass between $10^{3-5}$ $\rm{M_{\odot}}~{}$forms and collapses directly to a near-equivalent mass black hole. Such heavy seeds are regularly invoked to explain the existence of MBH at high-z, while the chemical signatures of the progenitor super massive star (SMS) phase may already have been detected (Nandal et al., 2024a, b; Schaerer et al., 2024). However, in this study, we will focus on light seeds and how rapid growth onto light seeds can provide an additional channel to MBH seeding.
@@
To understand the potential hurdles to the growth of light seeds, we study here whether rapid growth of light seeds can be achieved within gas-rich galaxies at high redshift. We conducted simulations using the Arepo code (Springel, 2010; Pakmor et al., 2016), incorporating self-consistent PopIII star formation, MBH formation, as well as SNe feedback. Our newly developed modules are designed to simulate the creation and accretion of sink particles, which serve as analogues to stars during their stellar lifetimes and subsequently evolve into BHs.
@@
The paper is structured as follows. In §2, we discuss the numerical methodology of our work. We briefly describe the working of the Arepo code. We also discuss in depth the initial conditions used and the sink particle techniques employed. §3 describes our results for two realisations (with and without SNe feedback). We compare our study with others from the literature in §4 and discuss some of our caveats in §5. Finally, we conclude our work in §6.
2 Numerical methods
To carry out the simulations described in this paper, we use the moving mesh code Arepo (Springel, 2010; Pakmor et al., 2016), which solves the Euler equations on an unstructured Voronoi tessellation that moves with the flow. The fluid dynamics are solved using a finite-volume, second-order reconstruction scheme coupled with an exact Riemann solver. The unstructured mesh adapts by locally refining and de-refining in a quasi-Lagrangian way to maintain a typical gas mass target per cell. We employ the hierarchical time integration scheme described in Springel et al. (2021), which splits the Hamiltonian into “fast” and “slow” components, i.e. those on short versus long time-steps, respectively. With this approach, interactions between particles/cells on the smallest time steps (up to a maximum of 3000 particles) are solved using direct summation, while a standard octree approach is used for all other interactions. This is primarily done for efficiency, to avoid unnecessary tree constructions for small numbers of particles, but has the added benefit of more accurately calculating the gravitational forces between black holes and their nearest neighbours, which are on the shortest time-steps, via direct summation (Bourne et al., 2024).
Within our suite of simulations, the dark matter (DM) particles have physical gravitational softening lengths $\Delta x$ = $10^{-2}$ pc, while gas cell softening adapt with the cell size down to a minimum of $\Delta x$ = $10^{-2}$ pc. This is the minimum cell length we allow and is controlled by the MinVolume parameter. Refinement is controlled via the JeansRefinement parameter with the number of cells used to resolve the Jeans length of the gas set to 8. We also employ a gas temperature floor of $\sim 3$ K.
@@ 2.1 Initial conditions
We create the initial conditions (ICs) for our idealised and isolated galaxies using GalIC (Yurin & Springel, 2014), which constructs galaxy models in collisionless equilibrium made up of both DM and other collisionless particles which make up the disk and bulge components. To accurately and realistically model early high redshift galaxies, we initialise our galaxies using the results from Chen et al. 2014, who found that for halo masses between $10^{7}$ $\rm{M_{\odot}}~{}$and $10^{8.5}$ $\rm{M_{\odot}}~{}$, the gas fraction to the total halo mass is $0.1$, lower than the cosmic baryon fraction. The reason for this is that in small early galaxies, SNe feedback is very effective in removing gas from galaxies, at least, until galaxies become more massive. As a result, we create our idealised galaxies with $90\%$ of the galaxy’s mass in the DM halo. The remaining mass of the galaxy being baryonic.
The DM halo is modelled as a NFW (Navarro et al., 1997) halo profile, while the gas is modelled as a spherical bulge with a radius $0.1$ times the radius of the parent halo. We set the virialized circular velocity of the halo to be $10$ km/s, which constrains the total mass of the galaxy to be $2.32\times 10^{8}$ $\ \text{M}_{\odot}$. We create the IC using $10^{6}$ DM particles and $10^{4}$ gas particles. The (initial) mass resolution is then $\sim 200$ $\rm{M_{\odot}}~{}$and $\sim 2000$ $\ \text{M}_{\odot}$ for the DM and gas particles, respectively. During the course of the simulation, the mass resolution of the gas cells refine based on local gas properties (specifically the Jeans length) and more gas cells are added to the grid as discussed in §2. The refinement causes the gas cell mass to range from approximately one to a few thousand solar masses. The DM particle resolution remains fixed throughout the simulation, but at $\sim 200$ $\rm{M_{\odot}}$, it is highly resolved. In addition to the DM particles and gas particles we also include $10^{4}$ collisionless ‘star’ particles which reside in a plane around the halo centre. These "stars" model the previous generation of stars inherited via the hierarchical build up of the galactic structure. The total mass in these genetic stars is $2.32\times 10^{6}$ and they add extra potential to the halo centre, helping to stabilise the inner part of the galaxy. The different components of the halo are outlined in Table 1.
@@
Following the creation of the initial conditions using GalIC, the output must be re-formatted into an output suitable for Arepo . To enable this, GalIC generated ICs must be inserted into our 20 kpc box and evolved using Arepo , this requires that the smoothed particle hydrodynamic (SPH) particles created by GalIC be converted into mesh cells. This is achieved using the ADD_BACKGROUND functionality (Springel, 2010; Weinberger et al., 2020) in Arepo . This procedure of taking the idealised halo and inserting it into an Arepo box inevitably results in some mass loss from the halo. This is because a small number of particles are scattered far from the host halo due to dynamical interactions at the outskirts of the halo. Therefore, these particles leave the 20 kpc box (isolated boundary conditions) in which we conduct the Arepo simulation. The final result is therefore a halo of mass $1.78\times 10^{8}$ $\ \text{M}_{\odot}$ in a box of $20$ kpc.